OpenStudy (anonymous):

Matilda invested $5,000 at a rate of 7.5% How much did she have after 6 months ? I GOT 23,750 BUT I AM NOT SURE IF THATS THE ANSWER CAN SOMEON HELP ME ?

OpenStudy (anonymous):

whats the equations again? i forgot it back in precalc lol

OpenStudy (anonymous):

I=PRT Principle * Rate * Time

OpenStudy (anonymous):

what are the units for time?

OpenStudy (anonymous):

days? months? weeks?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

oh lol nvm its prolly months but it could be years

OpenStudy (anonymous):

yea uhm do you noe if my answer is right ?

OpenStudy (anonymous):

r u a 5th grader

OpenStudy (anonymous):

oo a 8th grader but i forgot how to do this lmaoo (offensive )

OpenStudy (p0sitr0n):

5000* 1.075^6

OpenStudy (anonymous):

7.5% simple interest or compound interest?

OpenStudy (p0sitr0n):

depends if its calculated annually or 1/month

OpenStudy (anonymous):

i did 6/12 and i got o.5 and then with that i multipilyed 5000 * 7.5 * 0.5 and then i got 18750 +5000 and i got 23,750 do you think this is right ? oh and foolformath is simple interest

OpenStudy (anonymous):

it really depends on how often the interest is compounded i assumed monthly, because if it is compound annually you don't get any interest until the end of the year, so you would only have what you started with

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

If it is simple interest then the interest after 6 months is \( (5000 \times 7.5 \times1/2)/100= 187.5 \) So the total amount after 6 months is 5187.5

OpenStudy (p0sitr0n):

@FoolForMath isnt it supposed to be exponential? or we could do it both ways?

OpenStudy (anonymous):

but where did you get 1/2 from ?

OpenStudy (anonymous):

Simple interest is not exponential.

OpenStudy (anonymous):

6 months is half (1/2) year.

OpenStudy (p0sitr0n):

5000 = initial value 100% +7.5% = 1.075 6 = nb of months 5000* 1.075^6 = 7716.5

OpenStudy (anonymous):

and if its 3 month than i divide that by 12 ?

OpenStudy (anonymous):

@P0sitr0n: First I think that is compound interest and second it appears you are doing it for 6 years instead of 6 months.

OpenStudy (anonymous):

@eliiloveu: 12 months = 1 year 3 months = 3/12 = 1/4 year.

OpenStudy (anonymous):

ohh ok okay thanks can you ask me a question like the one i had on top to see if i know please if you have time because i have a test tommorow and i am scared if i get confused ?

OpenStudy (anonymous):

Okay, Matilda invested $5,00 at a rate of 7% How much did she have after 3 months ?

OpenStudy (anonymous):

8.75

OpenStudy (anonymous):

Right! that's the simple interest.

OpenStudy (anonymous):

oh thank you very much :)

OpenStudy (anonymous):

glad to help :)

OpenStudy (anonymous):

:)